OpenStudy (anonymous):

1. Graph the system of inequalities. Then use your graph to identify the point that represents a solution to the system. x > 3 y ≤ 2x – 5 A. (5, 6) B. (4, –1) C. (–3, –1) D. (1, 11) 2. Graph the system of inequalities. Then use your graph to identify the point that represents a solution to the system. x + y ≥ 5 x – 2y > 8 A. (6, 1) B. (6, –2) C. (8, –1) D. (6, 2)

OpenStudy (anonymous):

3. Graph the system of inequalities. Then use your graph to identify the point that represents a solution to the system. 3x – 4y > 0 x – 5y > 0 A. (6, 2) B. (–4, –2) C. (0, 0) D. (6, 0)

OpenStudy (anonymous):

Just graph the straight lines and then use the inequality to note where to shade!

OpenStudy (anonymous):

that don't tell me what (x, y) is

OpenStudy (anonymous):

it does, if the given point lies in the shaded region, then its the solution!

OpenStudy (anonymous):

i figured them out with a graphing calculator

OpenStudy (texaschic101):

you could have just plugged your answer choices into the inequalities and saw which satisfied the inequalities. For instance, the first one... x > 3 ---- (4,-1) 4 > 3 -- correct y <= 2x - 5 -1 <= 2(4) - 5 -1 <= 8 - 5 -1 <= 3 -- correct answer is (4,1) It is not that hard....just a little time consuming

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

but the question asks her to graph it.. !! u may lose points!

OpenStudy (texaschic101):

graphing will have to be done.

OpenStudy (anonymous):

the teacher doesn't see the graphing

OpenStudy (anonymous):

its online so she just shows us how to graph it

OpenStudy (texaschic101):

oh...really ? You might want to do the graphing anyways, just so you can learn how

OpenStudy (anonymous):

she shows us how I just forgot so I need to go back and look at how she did it

OpenStudy (texaschic101):

good idea :)

OpenStudy (anonymous):

thanks texaschick101

OpenStudy (texaschic101):

anytime :)